KyledaGreat:

Find a formula for the inverse of the following function, if possible. W(x)=4x^1/5+3 Selecting a radio button will replace the entered answer value(s) with the radio button value. If the radio button is not selected, the entered answer is used. W^-1(x) = ___________________ does not have an inverse function

KyledaGreat:

@tranquility

Tranquility:

To clarify, is the function: \( W(x) = \dfrac{4x^1}{5} + 3\)

KyledaGreat:

no

KyledaGreat:

\[W(x) = 4^{\frac{ 1 }{ 5}} + 3\]

Tranquility:

Oh \(\large W(x) = 4x^{\frac{1}{5}} + 3\)

KyledaGreat:

yes

KyledaGreat:

there's not a function ?

Tranquility:

When you're finding the inverse of a function, you have to switch the variables and solve for y. W(x) is equivalent to 'y' \(\large y = 4x ^{\frac{1}{5}} + 3\) Switch the x to a y, and the y to an x. And then solve for y

KyledaGreat:

i couldn't find a solution

Tranquility:

After you subtract 3 and divide by 4, you would have to ^5 both sides

KyledaGreat:

-243/1024

Tranquility:

You lost the x somewhere in the process \(\Large \frac{x-3}{4} = y^{\frac{1}{5}}\) and then you do it to the power of 5 on both sides to get rid of the power of 1/5

KyledaGreat:

1/5 = y^5

KyledaGreat:

right ?

Tranquility:

no? When you take the power of 5 on both sides, you'll get \(\Large (\frac{x-3}{4})^5= (y^{\frac{1}{5}})^5\) and \( ( y^{\frac{1}{5}} )^5 = y\)